# Open Your Ears to Take a Look: A State-of-the-Art Report on the Integration of Sonification and Visualization

### v1

@@ 5.3 Arts
During the classification of the corpus, an “art” category emerged, defining projects combining visualization and sonification for artistic purposes. We understand the emergence of art practices that intertwine sonification and visualization not only as inspiring for designers and audiences alike, but also relevant to the broader discourse around data representation towards a better, more efficient, and engaging human-data relationship.
@@
Working with the literature, we understood artistic endeavors call for a different kind of classification than idioms intended for audiovisual data analysis. We decided to not include such contributions in the systematic part of this STAR even when they complied with the requirement of being published in an academic journal and dealing with the simultaneous combination of visualization and sonification. We excluded papers where authors made their intentions of an artistic contribution explicit in the paper.
@@
Nevertheless, to give the reader a brief introduction to the field, we present representative cases of an emerging art category that could form the basis of future research on the integration of sonification and visualization. In addition to the artistic cases that emerged from our literature search, we use four more sources to curate a collection of representative cases for artistic endeavors: The Data Sonification Archive (DSA), the Computer Music Journal, the Leonardo Music Journal, and the Ars Electronica Festival. The Data Sonification Archive is an online crowd-sourced collection of data sonification projects. Out of 455 cases currently hosted in the DSA, three of the co-authors of this report recently categorized 139 cases as “art”, i.e., as having their primary purpose in creating and delivering an artistic experience to an audience [LLC23]. A preliminary search on some of the major venues for academic artistic publications – Computer Music Journal and Leonardo Music Journal – accounts for 182 and 118 cases respectively, in line with the volume of art-related cases in the Data Sonification Archive. A search of the archive of the renowned Ars Electronica Festival on projects that translate data into an audiovisual idiom returns 32 cases, again, in line with the 23 projects currently hosted on the DSA.
@@
Typically, data-driven artworks that use an audiovisual design address a non-expert audience and can take different forms such as live performances, multimedia installations, video productions, or web experiences where the sonification and visualization of the same dataset are combined. The DSA classification highlights recurring topics such as climate change and/or the representation of environmental data; individual and collective interaction in the urban space; internet and social media. The predominance of socially relevant topics shows that the emerging art category presents borderline characteristics with the adjacent category (in the DSA classification system, see [LLC23]) of “public engagement”. This latter category identifies projects that, although being often presented in the form of an artistic experience, have explicitly and primarily the goal of increasing awareness – even fostering activism – among the general public, on a specific phenomenon.
@@
This is the case for projects that present data related to climate change such as Klima|Anlage—Performing climate data [$*{}^{*}$19], one of the cases identified as artistic in our literature search. In this work, climate data from 1950 to 2100 can be chosen by the listener for twelve selected regions of the world and interactively sonified and visualized in an immersive installation environment with the goal of contributing to “the urgent need to inform the general public about climate change”.
@@
In Too Blue [Foo15], hosted on the DSA, the author created an audiovisual experience (in the form of a video) where music generated by tracking the land loss in coastal Louisiana over 78 years due to man-made levees, drilling and dredging for oil and gas, and climate change is combined with an aerial color-coded visual map that shows the increase in the “blues” colored areas, corresponding to the sea, in the coastal land of Louisiana. In Heat and the Heartbeat of the City [Pol06], the urgency to communicate climate change to the broader audience combines with a reflection on the urban space through the simultaneous sonification and visualization of actual and projected data on increased temperature in New York City. In aqua_forensic [SŠ18] data on invisible pollutants in the Danube river’s water, and the effect on its ecosystem, are collected, sonified, and visualized ‘to present the results of this scientific research to the wide audience in a poetic and artistic way’. Other prominent social issues such as personal data collection and privacy in the age of social media are at the center of multi-awarded art projects such as Digital Violence [FE21] where sonification is explicitly used to increase the audience’s engagement with the visualized dataset on the issue of illegal government surveillance practices. From the collective to the individual experience, projects such as Orbuculum [Ylm23] and Deep Sync [KOR23] invite the public to interactively engage with personal psychological and physiological dimensions to create an immersive auditory and visual experience. In a more-than-human perspective, works such as Spider Web Sonification [$*{}^{*}$20] and Biota Beats [$*{}^{*}$20] shift the focus from humans to other species in an effort to support the public engagement with science.
@@
Data art can stimulate an “artistic affectivization” that can contribute to building shared perspectives within a community [BMLA22]. The engagement of the audience with the phenomenon behind the data, however, also implies an information and knowledge transfer [$*{}^{*}$10] i.e. the public has not only to connect emotionally with the data, but also understand their meaning. The emergence of a category of artistic projects that combine sonification and visualization with the explicit intent to activate the audience on a socially relevant issue shows that multimodal approaches have the potential to generate impact, not only for education and awareness but also for stimulating action.
@@ 6 Concluding Discussion and Future Work
In this STAR, we have provided an overview of the field of audiovisual idiom design tailored towards data exploration and presentation. We have used a variety of perspectives to classify the existing literature. For each of these categorizations, we also provide a brief discussion or reflection. In section 5, we offered insights into three topics that are adjacent to our report and can be highly inspirational for future developments. In the following section, we take a step back, adopt a global viewpoint, and discuss the broader insights, challenges, and exciting future research opportunities.
@@
The culture of providing supplemental material: A thriving future for the field of audiovisual idiom design also depends on the communities’ culture when providing demos for their designs as supplemental material. Out of 57 articles in our surveyed literature, only 29 provided a demo that is still available today. Five provided a demo that is no longer available, and 23 provided no demo at all. Online repositories such as Zenodo, hosted by CERN, or long-term archives such as Phaidra, hosted by the University of Vienna, allow the upload of videos and even the assignment of DOIs. While making sound available in a paper used to be an explicit challenge to the sonification field, it is not the case anymore. An additional visual representation of the sound can also be included in the paper, using a spectrogram of the sound itself or an iconic representation of it.
@@
Evaluating audiovisual idioms: It is often challenging to conduct systematic, well-informed, and above all, properly performed and reported evaluations of audiovisual idioms. Especially, when the goal is to compare different studies. As mentioned earlier, Rönnberg and Forsell suggested the standardization of questionnaires tailored to the field in order to increase the comparability between different studies [RF22]. Establishing this comparability is particularly difficult because the two senses are used simultaneously in an audiovisual display, and the evaluation methods must adapt to this characteristic of the design. Specifically, the existing methods focus on several aspects. One is user performance, which studies quantitative metrics such as task completion times and precision of responses. Another is user experience, which focuses more on qualitative metrics such as user engagement, and memorability. From the data we collected in our STAR, we can draw high-level conclusions regarding the type of evaluations that have been used within the corpus of literature.
@@
Most of the papers in the systematic literature review used sonification to support the perception of the data visualization, some also aimed to provide additional information via the auditory channel complementing the visualization, and an additional few aimed for public engagement or facilitating multi-user analysis of the data. Among the papers that included a user evaluation, most reported a subjectively stated benefit of sonification in perceiving and interpreting the visualization. Some subjective ratings also suggested that the use of sonification was engaging, could provide emotional content, and was experienced as interesting or pleasant. Other papers also reported objective measurements of increased user performance when sonification and visualization were used in combination compared to visualization alone. This was shown in terms of higher scores or increased accuracy when sonification was used. However, this increase in accuracy was reported to be related to increased task completion times. Similar effects have also been shown in other studies not included in this systematic review [MBKSM16, Rön19a, Rön19b]. These results indicate that there is, in general, a benefit of combining visualization and sonification for data exploration and analysis, regardless of the specific visualization and/or sonification design. This benefit can be manifested in higher user engagement, increased user performance with higher scores or accuracy, or with more information conveyed to the user via both modalities. However, improvements in accuracy might come with the cost of increased task completion times. The case might be that the use of the additional information provided by sonification that makes the increased accuracy possible takes longer time to process and assess. While traditionally increased task completion times are considered a disadvantage, in some contexts this can be seen as positive. Slowing down the user resulting in them spending more time with their data, studying it from more and from different perspectives (including the auditory perspective), might reveal structures or patterns in a dataset, that would have stayed silent and unseen if analyzed with a conventional uni-modal display [HAS11].
@@
The potential of unconventional task distribution: The data from our STAR shows that reading levels [Ber83] are not equally distributed to the two senses: The vast majority – 50 out of 57 papers used the reading level “whole” for their visualization (see Table 2), while only about half of the papers use the same reading level for their sonification parts. This observation might be directly related to the way audiovisual idioms distribute the tasks of overview and detail to the two senses. The currently more popular distribution seems to be to use visualization to display an overview of the data while sonification is employed to provide details. While this distribution seems like an intuitive choice, future research should study the potential of switching the roles between sonification and visualization. It is our daily lives and the way we experience the world around us, that suggests to sonify an overview and to visualize details. Our auditory sense constantly screens our 360° environment while our visual sense actively focuses on details in our environment. The field of auditory process monitoring makes use of exactly this situation when alerting us of a status, regardless of our current visual focus. Nevertheless, we identified only three cases where the reading level for the sonification was “whole” and the one for the visualization was “group” [DLVDCG22, FBC12, YH18].
@@
A second phenomenon we observe in the data is that many audiovisual display idioms employ the same search level [Mun15] (see Figure 8) and most of them employ the search level of “explore” (see Table 2). Again, it seems like a fruitful future endeavor to study the opportunities of designs that explicitly break such a pattern.
@@
The potential of integrating sonification and scientific visualization: Most articles in our database can be regarded as related to the field of information visualization, rather than scientific visualization. Scientific visualization typically displays inherently spatial phenomena and is often used to represent data that varies over time. The inherently spatio-temporal quality of such displays constitutes their potential to be integrated with sonification, as sonification is a display technique tailored toward temporal data structures [$*{}^{*}$23]. We relate this argument back to the real-world example from our introduction where the combination of our visual and our auditory senses offer a better understanding of a phenomenon than by each sense alone: Rain falling outside of a closed window. Opening the window, to not only see but also hear the sound of the rain gives us a better estimation of the amount and intensity of the rainfall. Abstractly, in such a situation we perceive a spatio-temporal phenomenon. It is most plausible that adding sound to a time series of scientific visualization can increase its holistic interpretability in an ecologically valid manner [Neu04]. This phenomenon is shown in two publications in the database, in the work by MacDonald et al. [$*{}^{*}$18], and the work by Temor et al. [$*{}^{*}$21]. Similarly, these ideas also apply to animations in information visualization where temporal aspects of the data might be visualized and sonified.
@@
Design frameworks: When it comes to the possibilities to design audiovisual display idioms, most researchers, as well as most domain experts, will not be able to successfully develop their own designs. Interdisciplinary knowledge bridging visualization, sonification, interactive design, and human perception is necessary to design effective, engaging, and re-usable audiovisual display idioms. Therefore, we cannot expect a domain expert to be able to quickly draft a prototype the same way they may be able to do using established visualization-only frameworks such as matplotlib [Hun07].
@@
In our database, we identified several design frameworks tackling this challenge [PC19, PCB23, LF21, DLVDCG22, KLTW17, CWM21]. Their primary focus is the design of the sonification part of the idiom, which is why we cannot speak of truly balanced contributions to both senses. A promising endeavor to enable truly balanced designs, with both the visualization and the sonification being equally well-informed, might be to not design them using one, but two separate frameworks, each specialized for its purpose. In this regard, the recent work by Reinsch and Hermann [HR21, RH22, RH23] is promising, as it provides a sonification design framework embedded in the Python environment and is conceptually inspired by visualization design tools such as matplotlib. With both the sonification design and the visualization design happening in the same environment, in this case, for example, a Jupyter Notebook, both can be developed with the required quality to meet their individual standards. Regarding the introduction of users to a new design environment, be it for visualization, sonification, or their combination, effective “onboarding” is a pressing issue. In the visualization community, the topic has been studied in recent years [SPA23, $*{}^{*}$23], and the same will be necessary for sonification, and audiovisual display idioms. Again, when it comes to sonification as a technique to display data, the prior experience of domain experts or the general public is low. This lack of experience can, in the short term, only be met with carefully designed onboarding processes.
@@
Adjacent topics as inspiration to the field: Our brief look into the adjacent fields of monitoring, accessibility, and arts shows a special potential for inspiration for our communities’ future work. The field of monitoring offers established evaluation methods, that could potentially be adopted to support the evaluation of audiovisual idioms tailored towards exploratory data analysis. During our unstructured investigation into the field, we also identified the potential of a systematic STAR as a future contribution to the field of monitoring.
@@
Also, the adjacent field of accessibility can inspire future research, where understanding the analysis patterns of visually impaired individuals can inform design decisions for idioms tailored towards sighted people [WM10]. While the field of accessibility can be informative to the field of audiovisual audiovisual display idioms, the same holds true the other way around. Especially the designs from our STAR that use a redundant mapping to both the visual and the auditory sense, which can inspire designs for visually impaired users. Furthermore, such designs have the potential to foster successful collaborative data analysis involving both sighted and blind users.
@@
The identified artistic contributions show great potential to be inspirational to the field of audiovisual idiom design as well. In the future, effort should be made by the research community to develop a specific framework for the evaluation and analysis of audiovisual idioms at the crossing of art and public engagement to explore how specific design strategies (e.g., interactivity and embodiment, as used in many of the cases presented in subsection 5.3) can be combined to support both affect and sense-making. Again, a systematic STAR dedicated to the artistic perspective would be a highly timely contribution to the community.
@@
An audiovisual analytics community: With this state-of-the-art report, we hope to contribute to the future establishment of systematic research on audiovisual display idiom design. The current development concerning the studied co-author network and the temporal development of publication numbers in the field is promising. We hope to reach researchers from both the – today often disjunct – visualization and sonification communities, inspiring them for future interdisciplinary collaborations, and to open their ears to take a look.
@@
As supplemental material to this STAR, we provide:
@@
In addition, the corpus of relevant papers with tags, publication metadata and fulltext of open accesss work is available as a Zotero library at https://www.zotero.org/groups/integrationsonificationvisualization/items.

### v2

@@ 5.3 Arts
During the classification of the corpus, an “art” category emerged, defining projects combining visualization and sonification for artistic purposes. We understand the emergence of art practices that intertwine sonification and visualization not only as inspiring for designers and audiences alike but also relevant to the broader discourse around data representation towards a better, more efficient, and engaging human-data relationship.
@@
Working with the literature, we understood artistic endeavors call for a different kind of classification than idioms intended for audiovisual data analysis. We decided not to include such contributions in the systematic part of this STAR and excluded papers where authors made their intentions of an artistic contribution explicit in the paper.
@@
Nevertheless, to give the reader a brief introduction to the field, we present representative cases of an emerging art category that could form the basis of future research on the integration of sonification and visualization. In addition to the artistic cases that emerged from our literature search, we use four more sources to curate a collection of representative cases for artistic endeavors: The Data Sonification Archive (DSA), the Computer Music Journal, the Leonardo Music Journal, and the Ars Electronica Festival. The Data Sonification Archive is an online crowd-sourced collection of data sonification projects. Out of 455 cases currently hosted in the DSA, two of the co-authors of this report recently categorized 139 cases as “art”, i.e., as having their primary purpose in creating and delivering an artistic experience to an audience [LLC23]. A preliminary search on some of the major venues for academic artistic publications – Computer Music Journal and Leonardo Music Journal – accounts for 182 and 118 cases, respectively, in line with the volume of art-related cases in the Data Sonification Archive. A search of the archive of the renowned Ars Electronica Festival on projects that translate data into an audiovisual idiom returns 32 cases, again, in line with the 23 projects currently hosted on the DSA.
@@
Typically, data-driven artworks that use an audiovisual design address a non-expert audience and can take different forms, such as live performances, multimedia installations, video productions, or web experiences where the sonification and visualization of the same dataset are combined. The DSA classification highlights recurring topics such as climate change and/or the representation of environmental data, as well as individual and collective interaction in the urban space, internet, and social media. The predominance of socially relevant topics shows that the emerging art category presents borderline characteristics with the adjacent category (in the DSA classification system, see [LLC23]) of “public engagement”. This latter category identifies projects that, although being often presented in the form of an artistic experience, have explicitly and primarily the goal of increasing awareness – even fostering activism – among the general public on a specific phenomenon.
@@
This is the case for projects that present data related to climate change, such as Klima|Anlage—Performing climate data [GHJ19], one of the cases identified as artistic in our literature search. In this work, climate data and model predictions between 1950 and 2100 can be chosen by the listener for twelve selected regions of the world and interactively sonified and visualized in an immersive installation environment with the goal of contributing to “the urgent need to inform the general public about climate change”.
@@
In Too Blue [Foo15], hosted on the DSA, the author created an audiovisual experience (in the form of a video) where music generated by tracking the land loss in coastal Louisiana over 78 years due to man-made levees, drilling and dredging for oil and gas, and climate change is combined with an aerial color-coded visual map that shows the increase in the “blues” colored areas, corresponding to the sea, in the coastal land of Louisiana. In Heat and the Heartbeat of the City [Pol06], the urgency to communicate climate change to the broader audience combines with a reflection on the urban space through the simultaneous sonification and visualization of actual and projected data on increased temperature in New York City. In aqua_forensic [SŠ18] data on invisible pollutants in the Danube river’s water, and the effect on its ecosystem, are collected, sonified, and visualized ‘to present the results of this scientific research to the wide audience in a poetic and artistic way.’ Other prominent social issues such as personal data collection and privacy in the age of social media are at the center of multi-awarded art projects such as Digital Violence [FE21] where sonification is explicitly used to increase the audience’s engagement with the visualized dataset on the issue of illegal government surveillance practices. From the collective to the individual experience, projects such as Orbuculum [Ylm23], and Deep Sync [KOR23] invite the public to interactively engage with personal psychological and physiological dimensions to create an immersive auditory and visual experience. In a more-than-human perspective, works such as Spider Web Sonification [SQS20] and Biota Beats [KGS20] shift the focus from humans to other species in an effort to support the public engagement with science.
@@
Data art can stimulate an “artistic affectivization” that can contribute to building shared perspectives within a community [BMLA22]. The engagement of the audience with the phenomenon behind the data, however, also implies an information and knowledge transfer [MVC10], i.e., the public not only has to connect emotionally with the data but also has to understand their meaning. The emergence of a category of artistic projects that combine sonification and visualization with the explicit intent to activate the audience on a socially relevant issue shows that multimodal approaches have the potential to generate impact, not only for education and awareness but also for stimulating action.
@@ 6 Concluding Discussion and Future Work
In this STAR, we have provided an overview of the field of audiovisual idiom design tailored toward data exploration and presentation. We have used a variety of perspectives to classify the existing literature, and, in section 5, we offered insights into three topics that are adjacent to our report and can be highly inspirational for future developments. For our categorizations, we also provided brief discussions or reflections as part of the individual subsections 3.2 to 3.10.
Before we adopt a global viewpoint and discuss broader insights, challenges, and exciting future research opportunities, we now want to mention some of the key insights and results from previous sections. Overall, our corpus holds a quite diverse field of combined designs, employing visual idioms such as scatter plots, maps, and networks, as well as sonification techniques such as parameter mapping sonification, audification, and earcons.
The analysis of employed reading levels [Ber83] of the different designs revealed an imbalance between the two modalities. The visualization part of the identified audiovisual idioms mostly works at higher reading levels than their sonification counterparts. This phenomenon suggests that many designs provide an overview of the data using their visualization, while the sonification supports the analysis of details.
Such design decisions are also related to the level of redundancy an audiovisual idiom employs, and mixed approaches, which represent parts of the data redundantly and other parts in a complementary manner, seem to show special potential. Such a selective alignment of mapping between the sonification and visualization might assist us in perceiving an idiom as integrated rather than as two separate displays coexisting next to each other.
Furthermore, the option to use sonification as part of an idiom’s design can directly influence the visual view as well. Thinking of idioms with “multiple views,” one of the views could, for example, be covered by an “auditory view.” For visualization designers, such considerations are especially relevant when space is limited, as is the case, for example, on smartwatches [LDIC21].
@@
The culture of providing supplemental material: A thriving future for the field of audiovisual idiom design also depends on the communities’ culture when providing demos for their designs as supplemental material. Out of 57 articles in our surveyed literature, only 29 provided a demo that is still available today. Five provided a demo that is no longer available, and 23 provided no demo at all. Online repositories such as Zenodo, hosted by CERN, the platform OSF, maintained by the Center for Open Science, or long-term archives such as Phaidra, hosted by the University of Vienna, allow the upload of videos and even the assignment of DOIs. While making sound available in a paper used to be an explicit challenge to the sonification field, it is not the case anymore. An additional visual representation of the sound can also be included in the paper, using a spectrogram of the sound itself or an iconic representation of it.
@@
Evaluating audiovisual idioms: It is often challenging to conduct systematic, well-informed, and, above all, properly performed and reported evaluations of audiovisual idioms, especially when the goal is to compare different studies. As mentioned earlier, Rönnberg and Forsell suggested the standardization of questionnaires tailored to the field in order to increase the comparability between different studies [RF22]. Establishing this comparability is particularly difficult because the two senses are used simultaneously in an audiovisual display, and the evaluation methods must adapt to this characteristic of the design. Specifically, the existing methods focus on several aspects. One is user performance, which studies quantitative metrics such as task completion times and precision of responses. Another is user experience, which focuses more on qualitative metrics such as user engagement and memorability. From the data we collected in our STAR, we can draw high-level conclusions regarding the type of evaluations that have been used within the corpus of literature.
@@
Among the papers that included a user evaluation (21 of the 57 papers), most reported a subjectively stated benefit of sonification in perceiving and interpreting the visualization. Some subjective ratings and feedback also suggested that the use of sonification was engaging (six papers considered this [HK22, RS22, EEBR21, KAV21, Rön21, DCM18]), could provide emotional content (mainly two papers [RS22, Rön21]), was experienced as interesting (five papers [HK22, RS22, EEBR21, Rön21, DCM18]), as well as pleasant (four papers [RS22, EEBR21, Rön21, DCM18]), and in perceived improved quality of visual information (six papers [PFH22, EEBR21, KAV21, FN18, RJ16, BF12]). Other papers also reported subjective and objective measurements of improved interpretation and higher understanding of data when sonification and visualization were used in combination (16 papers [PFH22, CWM21, EEBR21, DCM18, FN18, GDAS18, YH18, BBV16, GRK16, NSC16, RJ16, PBM15, PBV14, ASH12, BF12, Bea11]). This was shown in terms of higher scores or increased accuracy when sonification was used. However, this increase in accuracy was reported to be related to increased task completion times (see, for example, in [RJ16]). Similar effects have also been shown in other studies not included in this systematic review [MBKSM16, Rön19a, Rön19b]. The case might be that the use of the additional information provided by sonification that makes the increased accuracy possible takes a longer time to process and assess. While traditionally increased task completion times are considered a disadvantage, in some contexts, this can be seen as positive. Slowing down the user, resulting in them spending more time with their data, studying it from more and from different perspectives (including the auditory perspective), might reveal structures or patterns in a dataset that would have stayed silent and unseen if analyzed with a conventional unimodal display [HAS11].
These results indicate that there is, in general, a benefit of combining visualization and sonification for data exploration and analysis, regardless of the specific visualization and/or sonification design. Nevertheless, we consider it a major task for the future research of our community to systematically study what kinds of benefits the integration of sonification and visualization can provide.
@@
The potential of unconventional task distribution: The data from our STAR shows that reading levels [Ber83] are not equally distributed to the two senses: The vast majority – 50 out of 57 papers used the reading level “whole” for their visualization (see Table 2), while only about half of the papers use the same reading level for their sonification parts. This observation might be directly related to the way audiovisual idioms distribute the tasks of overview and detail to the two senses. The currently more popular distribution seems to be to use visualization to display an overview of the data, while sonification is employed to provide details. While this distribution seems like an intuitive choice, future research should study the potential of switching the roles between sonification and visualization. It is our daily lives and the way we experience the world around us that suggests sonifying an overview and visualizing details. Our auditory sense constantly screens our 360° environment while our visual sense actively focuses on details in our environment. The field of auditory process monitoring makes use of exactly this situation when alerting us of a status, regardless of our current visual focus. Nevertheless, we identified only three cases where the reading level for the sonification was “whole” and the one for the visualization was “group” [DLVDCG22, FBC12, YH18].
@@
A second phenomenon we observe in the data is that many audiovisual display idioms employ the same search level [Mun15] (see Figure 8), and most of them employ the search level of “explore” (see Table 2). Again, it seems like a fruitful future endeavor to study the opportunities of designs that explicitly break such a pattern.
@@
The potential of integrating sonification and scientific visualization: Scientific visualization typically displays inherently spatial phenomena and is often used to represent data that varies over time. The inherently spatio-temporal quality of such displays constitutes their potential to be integrated with sonification, as sonification is a display technique tailored toward temporal data structures [GGB05, ERI23]. We relate this argument back to the real-world example from our introduction, where the combination of our visual and auditory senses offers a better understanding of a phenomenon than each sense alone: Rain falling outside of a closed window. Opening the window to not only see but also hear the sound of the rain gives us a better estimation of the amount and intensity of the rainfall. Abstractly, in such a situation, we perceive a spatio-temporal phenomenon. It is most plausible that adding sound to a time series of scientific visualization can increase its holistic interpretability in an ecologically valid manner [Neu04]. This phenomenon is shown in two publications in the database, in the work by MacDonald et al. [MNW18], and the work by Temor et al. [TMN21]. Similarly, these ideas also apply to animations in information visualization, where temporal aspects of the data might be visualized and sonified.
@@
Design frameworks: When it comes to the possibilities of designing audiovisual display idioms, most researchers, as well as most domain experts, will not be able to develop their own designs. Interdisciplinary knowledge bridging visualization, sonification, interactive design, and human perception is necessary to design effective, engaging, and re-usable audiovisual display idioms. Therefore, we cannot expect a domain expert to be able to quickly draft a prototype the same way they may be able to do using established visualization-only frameworks such as matplotlib [Hun07].
@@
In our database, we identified several design frameworks tackling this challenge [PC19, PCB23, LF21, DLVDCG22, KLTW17, CWM21]. Their primary focus is the design of the sonification part of the idiom, which is why we cannot speak of truly balanced contributions to both senses. A promising endeavor to enable truly balanced designs, with both the visualization and the sonification being equally well-informed, might be to not design them using one but two separate frameworks, each specialized for its purpose. In this regard, the recent work by Reinsch and Hermann [HR21, RH22, RH23] is promising, as it provides a sonification design framework embedded in the Python environment and is conceptually inspired by visualization design tools such as matplotlib. With both the sonification design and the visualization design happening in the same environment, in this case, for example, a Jupyter Notebook, both can be developed with the required quality to meet their individual standards. Regarding the introduction of users to a new design environment, be it for visualization, sonification, or their combination, effective “onboarding” is a pressing issue. In the visualization community, the topic has been studied in recent years [SPA23, SWG23], and the same will be necessary for sonification and audiovisual display idioms. Again, when it comes to sonification as a technique to display data, the prior experience of domain experts or the general public is low. This lack of experience can, in the short term, only be met with carefully designed onboarding processes.
@@
Adjacent topics as inspiration to the field: Our brief look into the adjacent fields of monitoring, accessibility, and arts shows a special potential for inspiration for our communities’ future work. The field of monitoring offers established evaluation methods that could potentially be adopted to support the evaluation of audiovisual idioms tailored toward exploratory data analysis. During our unstructured investigation into the field, we also identified the potential of a systematic STAR as a future contribution to the field of monitoring.
@@
Also, the adjacent field of accessibility can inspire future research, where understanding the analysis patterns of visually impaired individuals can inform design decisions for idioms tailored towards sighted people [WM10]. While the field of accessibility can be informative to the field of audiovisual idioms, the same holds true the other way around. Especially the designs from our STAR that use redundant mapping to both the visual and the auditory senses can inspire future designs for visually impaired users. Furthermore, such designs have the potential to foster successful collaborative data analysis involving both sighted and blind users.
@@
The identified artistic contributions show great potential to be inspirational to the field of audiovisual idiom design as well. In the future, effort should be made by the research community to develop a specific framework for the evaluation and analysis of audiovisual idioms at the crossing of art and public engagement. Such efforts need to explore how specific design strategies (e.g., interactivity and embodiment, as used in many of the cases presented in subsection 5.3) can be combined to support both affect and sense-making. Again, a systematic STAR dedicated to the artistic perspective would be a highly timely contribution to the community.
@@
An audiovisual analytics community: With this state-of-the-art report, we hope to contribute to the future establishment of systematic research on audiovisual display idiom design. The current development concerning the studied co-author network and the temporal development of publication numbers in the field is promising. We hope to reach researchers from both the visualization and sonification communities, inspiring them for future interdisciplinary collaborations, and to open their ears and take a look.
@@
As supplemental material to this STAR, we provide at https://phaidra.fhstp.ac.at/o:5541:
@@
In addition, the corpus of relevant papers with tags, publication metadata, and full text of open access work is available as a Zotero library at https://www.zotero.org/groups/integrationsonificationvisualization/items.